Two tiny objects with equal charges of 8.57 μC are placed at the two lower corners of a square with sides of 0.131 m, as shown. Find the electric field at point B, midway between the upper left and right corners. If the direction of the electric field is upward, enter a positive value. If the direction of the electric field is downward, enter a negative value.
Answer #1

It is a square if we find radius then

r= d/2= 0.131/2=0.065

diagonal=√[0.065² + 0.065²] = 0.091 m and cos θ = 0.065/0.091 = 0.71

The direction will be upwards perpendicular to the side

E= 2*k*q*cos/r^2 ( because here q1=q2)

E = [{2*9*(10⁹)*0.71*8.57*(10⁻⁶)}/(0.091²)] = 1.32*107 N-c